Grade 12Algebra

What is the Common factor of (13^7+11^7) and (13^5+11^5)

Riya singh
6 Years agoGrade 12

1 Answer

Arun
6 Years ago
For n odd, we have the algebraic identities 
x^n + y^n = (x + y)(x^(n-1) - x^(n-2)y + x^(n-3)y^2 + ... + x^2 y^(n-3) - xy^(n-2) + y^(n-1)). 
 
Therefore, 
 
13^7 + 11^7 = (13 + 11)(13^6 - 13^5*11 + 13^4*11^2 - 13^3*11^3 + 13^2*11^4 - 13*11^5 + 11^6) 
 
13^5 + 11^5 = (13 + 11)(13^4 - 13^3*11 + 13^2*11^2 - 13*11^3 + 11^4) 
 
and it follows that 24 is a common factor of (13^7 + 11^7) and (13^5 + 11^5).